OpenStudy (anonymous):

Help Please!!!!!! Use a sum or difference identity to write sin(x-pi/4) as a function of x alone? Will give medal!!!!!! Please

OpenStudy (jdoe0001):

hmmm so.... what's are the \(\bf sin\left( \cfrac{\pi }{4}\right)\quad and\quad cos\left( \cfrac{\pi }{4}\right)\) anyway?

OpenStudy (anonymous):

Sin (pi/4) = sqrt 2/2 and cos (pi/4)= is the same thing

OpenStudy (jdoe0001):

ok... let us use that then \(\bf sin\left( x- \cfrac{\pi }{4}\right) \\ \quad \\ sin({\color{brown}{ \alpha}} - {\color{blue}{ \beta}})=sin({\color{brown}{ \alpha}})cos({\color{blue}{ \beta}})- cos({\color{brown}{ \alpha}})sin({\color{blue}{ \beta}})\qquad thus \\ \quad \\ sin(x)\cdot cos\left( \frac{\pi }{4} \right)-cos(x)\cdot sin\left( \frac{\pi }{4} \right)\quad \begin{cases} cos\left( \frac{\pi }{4} \right)=\cfrac{\sqrt{2}}{2}\\ sin\left( \frac{\pi }{4} \right)=\cfrac{\sqrt{2}}{2} \end{cases}\quad thus \\ \quad \\ sin(x)\cdot \cfrac{\sqrt{2}}{2}-cos(x)\cdot \cfrac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

Okay, would you keep going after that?

OpenStudy (jdoe0001):

hmmm missing a - there.. . one sec

OpenStudy (jdoe0001):

\(\bf sin(x)\cdot \cfrac{\sqrt{2}}{2}-cos(x)\cdot \cfrac{\sqrt{2}}{2}\implies \cfrac{\sqrt{2}sin(x)}{2}-\cfrac{\sqrt{2}cos(x)}{2} \\ \quad \\ \cfrac{\sqrt{2}sin(x)-\sqrt{2}cos(x)}{2}\implies \cfrac{\sqrt{2}[sin(x)-cos(x)]}{2}\)

OpenStudy (anonymous):

Alright awesome thank you!!

OpenStudy (jdoe0001):

yw